# A Geometry Map-Based Propagation Model for Urban Channels

### v1

@@ Abstract
With the rapid deployment of 5G and future 6G networks, accurate modeling of urban radio propagation has become critical for system design and network planning. However, conventional statistical or empirical models fail to fully capture the influence of detailed geometric features in dense urban environments. In this paper, we propose a geometry map-based propagation model that directly extracts key parameters from a 3D geometry map and incorporates the Uniform Theory of Diffraction (UTD) to recursively compute multiple diffraction fields, thereby enabling accurate prediction of large-scale path loss and time-varying Doppler characteristics in urban channels. A significant buildings identification algorithm is developed to efficiently detect buildings that significantly affect signal propagation. The proposed model is validated using urban measurement data, showing excellent agreement with path loss in both LOS and NLOS conditions. In particular, for NLOS scenarios with complex diffraction mechanisms, it outperforms the 3GPP and simplified models, reducing the RMSE by 7.1 dB and 3.18 dB, respectively. Doppler analysis further demonstrates its accuracy in capturing time-varying propagation characteristics, confirming the scalability and generalization capability of the model in urban environments.
@@ I Introduction
In general, urban channels exhibit the following characteristics [lyu2019characterizing]: i) pronounced multipath effects: Reflectors such as buildings and walls generate a large number of discrete multipath components (MPCs), leading to significantly greater delay spread and angular spread compared to open areas. ii) strong time variability and non-stationarity: The movement of the transmitter (TX) and receiver (RX), together with the presence of dynamic scatterers in the environment, causes the channel characteristics to evolve rapidly over time. iii) non-uniform angular spreads: In urban canyon scenarios, the azimuthal spread is usually much larger than the elevation spread, while intersections introduce additional scattering paths. iv) remarkable environmental dependence: Channel characteristics are highly dependent on street geometry and building distribution, exhibiting strong site-specific correlations. Therefore, channel models capable of accurately capturing urban environmental features and faithfully reflecting the propagation mechanisms are essential.
@@
Map-based channel models represent an important category of site-specific propagation models that integrate environmental geometry into channel characterization [10742569]. Unlike stochastic models that rely on statistical fitting, map-based models utilize digital maps or 3D geometric databases to reconstruct the propagation environment, enabling more accurate modeling of path loss, MPCs, and spatial correlation. Representative standardization efforts, such as the 3GPP TR 38.901 hybrid map-based model [zhu20213gpp], incorporate map-based extensions to enhance realism in urban and indoor scenarios by combining deterministic ray-tracing with geometry-informed stochastic processes. Similarly, the METIS [unknown] and WINNER II [kyosti2007winner] frameworks adopt map-based modeling to support location-dependent channel generation and mobility-aware simulations. These developments highlight the increasing emphasis on geometry-aware modeling as wireless communication systems evolve toward higher frequencies and denser deployments. Despite these advances, most standardized map-based models still rely on deterministic simulation engines, where environmental information serves only as static input rather than as part of an inference process. Therefore, establishing a propagation model that directly extracts and utilizes environmental parameters from geometric maps and that derives channel characteristics based on physical propagation principles is essential. Such a model would provide a scalable and physically interpretable framework for accurately predicting radio channel behavior across diverse urban canyon environments [he2024wireless].
@@
Several measurements and models for map-based channels have been carried out [kyosti2017map, priebe2012calibrated, kanhere2023calibration, gan2018hybrid, pascual2018wireless, carton2016validation, wang2021ray, ying2025site, lim2017map, zhu2018novel, zhu2022map, grosse2016hybrid, alkhateeb2019deepmimo, alrabeiah2020deep, alkhateeb2023deepsense, zhao2023nerf2, bakirtzis2022deepray, hoydis2023sionna, cao2024raypronet, qiu2022pseudo, chatelier2024model, fu2025ckmdiff]. These models mainly fall into three categories: i) geometry-based deterministic channel models (GDCMs): These models characterize MPC parameters in a fully deterministic manner, typically relying on ray-tracing methods [kanhere2024calibration]. Their accuracy depends on detailed knowledge of building layouts, street geometries, and material properties [10742569]. However, the strong reliance on environmental information limits their applicability in large-scale simulations, real-time evaluations, or scenarios with incomplete environmental data [rainer2021scalable]. ii) Hybrid map-based model: These models combine deterministic methods for dominant paths with statistical models for small-scale scattering [zhu2022map]-[grosse2016hybrid], balancing realism and efficiency. However, their parameters require extensive calibration, limiting accuracy and generalization with scarce data, and their adaptability in dynamic environments remains limited. iii) Data-driven approach: These models use high-resolution maps or point clouds to train neural networks for predicting path loss, delay–angle spectra, or radio maps, showing strong potential for rapid inference. Physics-informed learning and domain adaptation further reduce the gap between simulated and measured data [alkhateeb2019deepmimo, alrabeiah2020deep, alkhateeb2023deepsense, zhao2023nerf2, bakirtzis2022deepray, hoydis2023sionna, cao2024raypronet, qiu2022pseudo, chatelier2024model, fu2025ckmdiff]. However, their main limitations lie in the reliance on large amounts of labeled data, challenges in generalization, and the high retraining cost when the propagation environment undergoes significant changes.
@@ V CONCLUSION
In this paper, we propose a geometry map-based propagation model which directly extracts key geometric parameters from an urban 3D map database and, by incorporating the UTD, derives a recursive formulation for the calculation of multiple diffraction fields, enabling the derivation of essential channel characteristics such as path loss, multipath delay, and Doppler distributions. In addition, a significant buildings identification algorithm is developed to efficiently filter out building structures that have a significant impact on propagation paths. The model is validated based on measurements. The results show that in typical urban environments, the proposed model accurately characterizes the path loss variations under both LOS and NLOS conditions. Particularly in NLOS scenarios, where propagation mechanisms are more complex, the model achieves markedly higher prediction accuracy than the 3GPP standard model and simplified models neglecting multiple diffraction effects, reducing the RMSE by approximately 7.1 dB and 3.18 dB, respectively. Furthermore, Doppler spread analysis confirms that the model precisely characterizes the time-varying propagation behavior, achieving the closest agreement with measurements among all compared models, thereby validating scalability and generalization ability of the proposed model.

### v2

# A Geometry Map-Based Site-Specific Propagation Channel Model for Urban Scenarios

@@ Abstract
With the rapid deployments of 5G and 6G networks, accurate modeling of urban radio propagation has become critical for system design and network planning. However, conventional statistical or empirical models fail to fully capture the influence of detailed geometric features on site-specific channel variances in dense urban environments. In this paper, we propose a geometry map-based propagation channel model that directly extracts key parameters from a 3D geometry map and incorporates the Uniform Theory of Diffraction (UTD) to recursively compute multiple diffraction fields, thereby enabling accurate prediction of site-specific large-scale path loss and time-varying Doppler characteristics in urban scenarios. A well-designed identification algorithm is developed to efficiently detect buildings that significantly affect signal propagation. The proposed model is validated using urban measurement data, showing excellent agreement of path loss in both line-of-sight (LOS) and nonline-of-sight (NLOS) conditions. In particular, for NLOS scenarios with complex diffractions, it outperforms the 3GPP and simplified models, reducing the RMSE by 7.1 dB and 3.18 dB, respectively. Doppler analysis further demonstrates its accuracy in capturing time-varying propagation characteristics, confirming the scalability and generalization of the model in urban environments.
@@ I Introduction
In general, urban channels exhibit the following characteristics [lyu2019characterizing]: i) pronounced multipath effects: Reflectors such as buildings and walls generate a large number of discrete multipath components (MPCs), leading to significantly greater delay spread and angular spread compared to open areas. ii) strong time variability and non-stationarity: The movement of the transmitter (TX) and receiver (RX), together with the presence of dynamic scatterers in the environment, causes the channel characteristics to evolve rapidly over time. iii) non-uniform angular spreads: In urban canyon scenarios, the azimuthal spread is usually much larger than the elevation spread, while intersections introduce additional scattering paths. iv) remarkable environmental dependence: Channel characteristics are highly dependent on street geometry and building distribution, exhibiting strong site-specific correlations. Therefore, channel models capable of accurately capturing urban environmental features and faithfully reflecting the propagation mechanisms are essential [6805659].
@@
Map-based channel models represent an important category of site-specific propagation models that integrate environmental geometry into channel characterization [10742569]. Unlike stochastic models that rely on statistical fitting, map-based models utilize digital maps or 3D geometric databases to reconstruct the propagation environment, enabling more accurate modeling of path loss, MPCs, and spatial correlation. Representative standardization efforts, such as the 3GPP TR 38.901 hybrid map-based model [zhu20213gpp], incorporate map-based extensions to enhance realism in urban and indoor scenarios by combining deterministic ray-tracing with geometry-informed stochastic processes. Similarly, the METIS [unknown] and WINNER II [kyosti2007winner] frameworks adopt map-based modeling to support location-dependent channel generation and mobility-aware simulations. These developments highlight the increasing emphasis on geometry-aware modeling as wireless communication systems evolve toward higher frequencies and denser deployments. Despite these advances, most standardized map-based models still rely on deterministic simulation engines, where environmental information serves only as static input rather than as part of an inference process [10839242]. Therefore, establishing a propagation model that directly extracts and utilizes environmental parameters from geometric maps and that derives channel characteristics based on physical propagation principles is essential. Such a model would provide a scalable and physically interpretable framework for accurately predicting radio channel behavior across diverse urban canyon environments [he2024wireless].
@@
Several measurements and models for map-based channels have been carried out [kyosti2017map, priebe2012calibrated, kanhere2023calibration, gan2018hybrid, pascual2018wireless, carton2016validation, 10872967, wang2021ray, ying2025site, lim2017map, zhu2018novel, zhu2022map, grosse2016hybrid, alkhateeb2019deepmimo, alrabeiah2020deep, alkhateeb2023deepsense, zhao2023nerf2, bakirtzis2022deepray, hoydis2023sionna, 10238401, cao2024raypronet, qiu2022pseudo, chatelier2024model, fu2025ckmdiff]. These models mainly fall into three categories: i) geometry-based deterministic channel models (GDCMs): These models characterize MPC parameters in a fully deterministic manner, typically relying on ray-tracing methods [kanhere2024calibration]. Their accuracy depends on detailed knowledge of building layouts, street geometries, and material properties [10742569]. However, the strong reliance on environmental information limits their applicability in large-scale simulations, real-time evaluations, or scenarios with incomplete environmental data [rainer2021scalable]. ii) Hybrid map-based model: These models combine deterministic methods for dominant paths with statistical models for small-scale scattering [zhu2022map]-[grosse2016hybrid], balancing realism and efficiency. However, their parameters require extensive calibration, limiting accuracy and generalization with scarce data, and their adaptability in dynamic environments remains limited. iii) Data-driven approach: These models use high-resolution maps or point clouds to train neural networks for predicting path loss, delay–angle spectra, or radio maps, showing strong potential for rapid inference. Physics-informed learning and domain adaptation further reduce the gap between simulated and measured data [alkhateeb2019deepmimo, alrabeiah2020deep, alkhateeb2023deepsense, zhao2023nerf2, bakirtzis2022deepray, hoydis2023sionna, cao2024raypronet, qiu2022pseudo, chatelier2024model, fu2025ckmdiff, 10238401]. However, their main limitations lie in the reliance on large amounts of labeled data, challenges in generalization, and the high retraining cost when the propagation environment undergoes significant changes.
@@ V CONCLUSION
In this paper, we propose a geometry map-based site-specific propagation model which directly extracts key geometric parameters from an urban 3D map database and, by incorporating the UTD, derives a recursive formulation for the calculation of multiple diffraction fields, enabling the derivation of essential channel characteristics such as path loss, multipath delay, and Doppler distributions. In addition, a identification algorithm is developed to efficiently filter out building structures that have a significant impact on propagation paths. The model is validated based on realistic channel measurements. The results show that in typical urban environments, the proposed model accurately characterizes path loss variations under both LOS and NLOS conditions. Particularly in NLOS scenarios, the model achieves markedly higher prediction accuracy than the 3GPP model and simplified models neglecting multiple diffraction effects, reducing RMSE by approximately 7.1 dB and 3.18 dB, respectively. Furthermore, Doppler spread analysis confirms that the model precisely characterizes time-varying propagation behaviors, achieving the closest agreement with measurements among all the compared models, thereby validating scalability and generalization of the proposed model.